# CLoMo: Counterfactual Logical Modification with Large Language Models

### 29 Nov 2023 v1

@@ Abstract
In this study, we delve into the realm of counterfactual reasoning capabilities of large language models (LLMs). Our primary objective is to cultivate the counterfactual thought processes within LLMs and rigorously assess these processes for their validity. Specifically, we introduce a novel task, Counterfactual Logical Modification (CLoMo), and a high-quality human-annotated benchmark. In this task, LLMs must adeptly alter a given argumentative text to uphold a predetermined logical relationship. To effectively evaluate a generation model’s counterfactual capabilities, we propose an innovative evaluation metric, the Logic-Aware Counterfactual Score to directly evaluate the natural language output of LLMs instead of modeling the task as a multiple-choice problem. Analysis shows that the proposed automatic metric aligns well with human preference. Our experimental results show that while LLMs demonstrate a notable capacity for logical counterfactual thinking, there remains a discernible gap between their current abilities and human performance.
@@ 1 Introduction
Despite large language models Arkoudas (2023); OpenAI (2022) perform strikingly in plenty of reasoning benchmarks Cobbe et al. (2021); Hendrycks et al. (2021a), late studies observe internal inconsistency in their reasoning processes Saparov and He (2023); Arkoudas (2023). The inconsistency is attributed to misunderstanding and misapplication of logical relations. However, logical relations in complex language reasoning are not yet properly quantified and evaluated.
@@
We summarize the contributions of this paper as follows:
1. We propose the task of Counterfactual Logical Modification, and contribute a corresponding CLoMo. to evaluate the counterfactual reasoning capability of LLMs in the scenario of complicated textual logical reasoning.
2. We propose the Logic-Aware Counterfactual Score for the logically consistent generation of large language models.
3. We conduct experiments on representative LLMs (GPT-3.5, GPT-4) and smaller language models (the LLaMA and LLaMA 2 families) and find that CLoMo is a very challenging task and the counterfactual logical reasoning ability of the existing model needs to be improved.
@@ 2 Related Works
Complex reasoning has been highly concerned as a significant yet challenging task for inspecting advanced artificial intelligence. For example, for solving commonsense reasoning problems Talmor et al. (2019, 2021); Huang et al. (2019); Bhagavatula et al. (2020); Sap et al. (2019), the models Yasunaga et al. (2021, 2022); Liu et al. (2021); Huang et al. (2021b) are required to reasonable applying commonsense knowledge to conduct the reasoning process for the final answer. Furthermore, multi-step reasoning needs the models to perform multiple reasoning steps while maintaining consistency and faithfulness. To achieve this, synthetic compositional reasoning tasks Betz (2020); Tafjord et al. (2021); Han et al. (2022); Saparov and He (2023) incorporates first-order logic to inspect and improve models logical consistency Pan et al. (2023); Olausson et al. (2023); Sanyal et al. (2022); Yang and Deng (2021). Moreover, real-scenario compositional reasoning Yu et al. (2020); Liu et al. (2020); Dalvi et al. (2021); Huang et al. (2022) joins commonsense, consider the uncertainty of events in multi-step logical reasoning, which challenge current models Bao et al. (2023); Xu et al. (2023); Jiao et al. (2023, 2022); Huang et al. (2023a, 2021a) to solve real-world reasoning problems with faithfulness. Additionally, the more faithful models should be able to consider counterfactuals. For example, answering questions given counterfactual conditions Yu et al. (2023); Tandon et al. (2019); Qin et al. (2021), or narrating a counterfactual scenario Qin et al. (2019). However, previous studies on counterfactual reasoning barely pay attention to the logical consistency or faithfulness of models. Therefore, in this work, we propose the task of Counterfactual Logical Modification that challenges the model to satisfy a given logical relation restriction while generating counterfactuals.
@@ 3.2 Benchmark Construction
We then post-process the question and the annotation so that for each data point, Argumnet comes from the original context, Premise1 comes from the correct option, and Premise2 and ${}^{\prime}$ come from the annotation.
@@ 4 SES: Self-Evaluation Scores
We aim to use the proven complex reasoning capabilities of the large language model itself to perform faster and more efficient reasoning evaluations of complex reasoning tasks that do not have easy access to standard/human-tested answers. Specifically, we split the scenario of complex logical reasoning evaluation into several discriminative tasks that LLMs have already seen and have been heavily trained on through logical conceptual graphs in order for the model to perform high-precision reasoning evaluation. We then collect the evaluations of these simple tasks and compute the ratings of the complex reasoning tasks based on the structure of the logical concept graph.
@@
$c(r|p_{1},a)$
You are an expert in logic. <definition of relation>. In the following, you are given an Argument and a Premise. Is the Premise <relation> the Argument? Please think step by step, and then answer “yes” or “no”.
@@
$c(r|p_{2},a^{\prime})$
You are an expert in logic. <definition of relation>. In the following, you are given an Argument and a Premise. Is the Premise <relation> the Argument? Please think step by step, and then answer “yes” or “no”.
@@
$c(r|p_{2},a)$
You are an expert in logic. <definition of relation>. In the following, you are given an Argument and a Premise. Is the Premise <relation> the Argument? Please think step by step, and then answer “yes” or “no”.
@@
Decomposed Self-Evaluation Metric As the demonstrated complex logical reasoning ability of current large language models, we let the large language model estimate $\text{Pr}(r|p_{1},a)$, $\text{Pr}(r|p_{2},a^{\prime})$, and $\text{Pr}(r|p_{2},a)$, respectively. Specifically, we design a few binary classification tasks for the large language models, so that the probabilities are simplified to $c(r|p_{1},a)$, $c(r|p_{2},a^{\prime})$, and $c(r|p_{2},a)$$\in\{0,1\}$. Finally, the overall logical modification score is computed by:
SES NA SA S W
@@
Few
Zero
CoT
Inference-Only LLaMA-7b
0.340
0.435
0.440
LLaMA-13b
0.435
0.410
0.425
LLaMA 2-7b
0.270
0.145
0.115
LLaMA 2-13b
0.300
0.045
0.430
LoRA LLaMA-7b
0.260
0.285
0.295
LLaMA-13b
0.415
0.420
0.395
LLaMA 2-7b
0.085
0.135
0.145
LLaMA 2-13b
0.435
0.305
0.300
Fine-Tuning LLaMA-7b
0.385
0.445
0.085
LLaMA-13b
0.494
0.291
0.430
LLaMA 2-7b
0.633
0.557
0.494
LLaMA 2-13b
0.506
0.468
0.494
Table 4: SES results of LLaMA and LLaMA 2 families. Few: few-shot demonstrations. Zero: zero-shot demonstrations. CoT: chain-of-thought prompting.
Figure 3: Per-relation performances of LLaMA and LLaMA 2 families.
@@ 5.1 Main Results
We first evaluate the large language models. We evaluate GPT-3.5-Turbo and GPT-4, respectively. At the same time, we recruited experts in linguistics to annotate the answers with humans. Table 3 demonstrates the results.
@@
5.2 Results on Smaller Models
We further evaluate LLaMA and LLaMA 2 on CLoMo. We conduct fine-tuning with LoRA and full fine-tuning. We also directly evaluate the pre-trained models. We use three formats of training and testing data points, including few-shot demonstrations (Few), zero-shot demonstrations (Zero), and chain-of-thought prompting (CoT). The evaluation results are shown in Table 4.
In general, the pre-trained LLaMA and LLaMA 2 models achieve a certain level of counterfactual reasoning, and model scale is beneficial, but still fall short of large language models such as GPT-4. Fine-tuning with LoRA does not improve the performances. Full fine-tuning is more beneficial than fine-tuning with LoRA for this task. On the other hand, few-shot demonstrations in the data points help solve the task better. This suggests that the models can learn from imitation to solve this challenging task. Also, chain-of-thought prompting helps the counterfactual reasoning of the model to some extent.
In conclusion, smaller models in counterfactual logical reasoning still need to be improved.
5.3 Impacts of Logical Relations
Among the four logical relations, many of the models can not predict the Sufficient Assumption. Especially with zero-shot demonstration data. For example, both LLaMA 2-7b and LLaMA 2-13b fine-tuned with LoRA gets SES=0 in this logical relation. Therefore, Sufficient Assumption seems to be the hardest relation for the models. In contrast, the models generally perform better in Necessary Assumptions. Some models work better after fine-tuning. For example, LLaMA-7b improves from 0.4 to 0.429 after fine-tuned with LoRA and to 0.514 after full fine-tuning, and improves from 0.086 to 0.486 after fine-tuned with LoRA and 0.500 after full fine-tuning. While in some settings, for example, using the chain-of-thought prompting data, LLaMA-7b performs well with the pre-trained parameters, and fine-tuning with LoRA or full parameters both induce performance drop.
In conclusion, the extent to which the model understands different logical relations varies considerably. Therefore, an improved understanding of all logical relations is needed to improve models’ counterfactual logical reasoning ability.
@@ 5.3 Impacts of Logical Relations
Figure 4: A successful case of counterfactual modification generated by GPT-4, which makes a logically consistent ${}^{\prime}$ and is accordant with human reasoning.
@@
Figure 5: An inferior case of counterfactual modification generated by GPT-4. GPT-4 modifies Statement 1 by replacing intellectual activities with physical activities, and the logical relation restriction is not satisfied.
@@ 5.4 Case Study
We compare the modification by GPT-4 and humans.
Figure 4 shows a case about sufficient assumption. That is, whether premise is a sufficient assumption for argument. The focus of the two premises here is to predict the impact of the charter provision. Switching from Premise 1 to Premise 2, the focus of discussion changes from the group of major powers to other nations in response to threats to world peace. This mainly affects the elaboration of Statement 3. GPT-4 has revised Statement 3 accordingly. Humans also made changes to Statement 3. Both revisions emphasized the influence of the five major powers in Statement 3, which corresponded to Premise 2. We find that GPT-4 can handle complex logical reasoning and counterfactual reasoning to a certain extent.
Figure 5 is about weakening an argument. The argument is on human intellectual development. Statement 2 provides evidence to support Statement 1. Premise 1 is on medical conditions and treatments providing counterexamples, and Premise 2 is on inaccuracies in research data. Human modifies Statement 2 to emphasize conversation, thus satisfying the logical conflict with the inaccuracies in research data described in Premise 2, thus satisfying the weaken relation. GPT-4 modifies Statement 1 by replacing intellectual activities with physical activities such as sports or gym. Intellectual/physical activities have some counterfactual contrast, but in the context of the argument at hand, the logical relation restriction is not satisfied. We find that GPT-4 can still be flawed in counterfactual reasoning.
In conclusion, complex counterfactual reasoning is challenging for large language models and needs improvements.

### 14 Feb 2024 v3

@@ Abstract
In this study, we delve into the realm of counterfactual reasoning capabilities of large language models (LLMs). Our primary objective is to cultivate the counterfactual thought processes within LLMs and rigorously assess these processes for their validity. Specifically, we introduce a novel task, Counterfactual Logical Modification (CLoMo), and a high-quality human-annotated benchmark. In this task, LLMs must adeptly alter a given argumentative text to uphold a predetermined logical relationship. To effectively evaluate a generation model’s counterfactual capabilities, we propose an innovative evaluation metric, the decomposed Self-Evaluation Score (SES) to directly evaluate the natural language output of LLMs instead of modeling the task as a multiple-choice problem. Analysis shows that the proposed automatic metric aligns well with human preference. Our experimental results show that while LLMs demonstrate a notable capacity for logical counterfactual thinking, there remains a discernible gap between their current abilities and human performance. We will release the data and code after publication.
@@ 1 Introduction
Despite large language models Arkoudas (2023); OpenAI (2022) perform strikingly in plenty of reasoning benchmarks Cobbe et al. (2021); Hendrycks et al. (2021a), late studies observe an internal inconsistency in their reasoning processes Saparov and He (2023); Arkoudas (2023). The inconsistency is attributed to misunderstanding and misapplication of logical relations. However, logical relations in complex language reasoning are not yet properly quantified and evaluated.
@@
The contributions of this paper are three-fold:
  • We propose the task of Counterfactual Logical Modification and contribute a corresponding CLoMo. to evaluate the counterfactual reasoning capability of LLMs in the scenario of complicated textual logical reasoning.
  • We propose the decomposed Self-Evaluation Score (SES) for the logically consistent generation of large language models.
  • We conduct experiments on representative LLMs (GPT-3.5, GPT-4) and small language models (the LLaMA and LLaMA 2 families) and find that CLoMo is a very challenging task and the counterfactual logical reasoning ability of the existing model needs to be improved.
@@ 2 Related Works
Complex reasoning has been highly concerned as a significant yet challenging task for inspecting advanced artificial intelligence. For example, for solving commonsense reasoning problems Talmor et al. (2019, 2021); Huang et al. (2019); Bhagavatula et al. (2020); Sap et al. (2019), the models Yasunaga et al. (2021, 2022); Liu et al. (2021); Huang et al. (2021b) are required to reasonable applying commonsense knowledge to conduct the reasoning process for the final answer. Furthermore, multi-step reasoning needs the models to perform multiple reasoning steps while maintaining consistency and faithfulness. To achieve this, synthetic compositional reasoning tasks Betz (2020); Tafjord et al. (2021); Han et al. (2022); Saparov and He (2023) incorporate first-order logic to inspect and improve models logical consistency Pan et al. (2023); Olausson et al. (2023); Sanyal et al. (2022); Yang and Deng (2021). Moreover, real-scenario compositional reasoning Yu et al. (2020); Liu et al. (2020); Dalvi et al. (2021); Huang et al. (2022) joins commonsense, consider the uncertainty of events in multi-step logical reasoning, which challenge current models Bao et al. (2023); Xu et al. (2023); Jiao et al. (2023, 2022); Huang et al. (2023a, 2021a) to solve real-world reasoning problems with faithfulness. Additionally, the more faithful models should be able to consider counterfactuals. For example, answering questions given counterfactual conditions Yu et al. (2023); Tandon et al. (2019); Qin et al. (2021), or narrating a counterfactual scenario Qin et al. (2019). However, previous studies on counterfactual reasoning barely pay attention to the logical consistency or faithfulness of models. Therefore, in this work, we propose the task of Counterfactual Logical Modification that challenges the model to satisfy a given logical relation restriction while generating counterfactuals.
@@ 3.2 Benchmark Construction
We then post-process the question and the annotation so that for each data point, Argument comes from the original context, Premise1 comes from the correct option, and Premise2 and ${}^{\prime}$ come from the annotation.
@@ 4 SES: Self-Evaluation Scores
We aim to use the proven complex reasoning capabilities of the large language model itself to perform faster and more efficient reasoning evaluations of complex reasoning tasks that do not have easy access to standard/human-tested answers. Specifically, we split the scenario of complex logical reasoning evaluation into several discriminative tasks that LLMs have already seen and have been heavily trained on through logical conceptual graphs for the model to perform high-precision reasoning evaluation. We then collect the evaluations of these simple tasks and compute the ratings of the complex reasoning tasks based on the structure of the logical concept graph.
@@
$c(r|p_{1},a)$ You are an expert in logic. <definition of relation>. In the following, you are given an Argument and a Premise. Is the Premise <relation> the Argument? Please think step by step, and then answer ‘‘yes’’ or ‘‘no’’.
@@
$c(r|p_{2},a^{\prime})$ You are an expert in logic. <definition of relation>. In the following, you are given an Argument and a Premise. Is the Premise <relation> the Argument? Please think step by step, and then answer ‘‘yes’’ or ‘‘no’’.
@@
$c(r|p_{2},a)$ You are an expert in logic. <definition of relation>. In the following, you are given an Argument and a Premise. Is the Premise <relation> the Argument? Please think step by step, and then answer ‘‘yes’’ or ‘‘no’’.
@@
[…]
@@
Decomposed Self-Evaluation Score As the demonstrated complex logical reasoning ability of current large language models, we let the large language model estimate $\text{Pr}(r|p_{1},a)$, $\text{Pr}(r|p_{2},a^{\prime})$, and $\text{Pr}(r|p_{2},a)$, respectively. Specifically, we design a few binary classification tasks for the large language models, so that the probabilities are simplified to $c(r|p_{1},a)$, $c(r|p_{2},a^{\prime})$, and $c(r|p_{2},a)$$\in\{0,1\}$. Finally, the overall logical modification score is computed by:
Figure 3: Per-relation performances of LLaMA and LLaMA 2 families. NA: Necessary Assumption. SA: Sufficient Assumption. S: Strengthen. W: Weaken.
@@ 5.1 Main Results
We first evaluate the large language models GPT-3.5-Turbo OpenAI (2022) and GPT-4 OpenAI (2023). We also recruit 2 domain experts to contribute to the human performance. Table 3 demonstrates the results.
@@
5.2 Fine-Tuning with Counterfactual Data
We then fine-tune LLaMA Touvron et al. (2023a) and LLaMA 2 Touvron et al. (2023b). We randomly split the CLoMo dataset into 60%/20%/20% training/development/test data, and fine-tune the models with the CLoMo training data. The implementation details are explained in Appendix C. We also compare the results with inference-only settings. The evaluation results are demonstrated in Figure 3.
In general, the pre-trained LLaMA and LLaMA 2 models achieve a certain level of counterfactual reasoning, and fine-tuning with counterfactual data further improves the performance. For example, LLaMA 2-13B with chain-of-thought prompting increases by 38.4% (SES: 0.430 $\to$ 0.595), with few-shot prompting increases by 98.3% (SES: 0.300 $\to$ 0.595), and with zero-shot prompting increases by 1,188.9% (SES: 0.045 $\to$ 0.580). The results indicate that such counterfactual data are barely seen in the LLaMA and LLaMA 2 pre-training data, and datasets such as the proposed CLoMo are much needed for developing models’ counterfactual reasoning capabilities.
Moreover, among the four logical relations, Weaken performances are significantly increased after fine-tuning. However, the absolute SES scores of Sufficient Assumption problems after fine-tuning are still relatively low. It shows that Sufficient Assumption are challenging. Therefore, we still need a profound investigation of different logical relations to improve the models’ counterfactual logical reasoning ability.
[…]
CoT Flan-T5-Large Chung et al. (2022) 780M 0.28 0.30 0.21 0.46 0.18
Flan-T5-XL Chung et al. (2022) 3B 0.21 0.18 0.07 0.34 0.19
Flan-T5-XXL Chung et al. (2022) 11B 0.26 0.33 0.07 0.37 0.17
ChatGLM2-6B Du et al. (2022a) 6B 0.25 0.32 0.07 0.34 0.17
Baichuan2-7B-Chat Baichuan (2023) 7B 0.24 0.20 0.07 0.49 0.18
Baichuan2-13B-Chat Baichuan (2023) 13B 0.26 0.19 0.21 0.49 0.22
InternLM-Chat-7B Team (2023) 7B 0.31 0.32 0.21 0.43 0.25
Vicuna-7B-v1.5 Chiang et al. (2023) 7B 0.15 0.13 0.00 0.29 0.14
Vicuna-13B-v1.5 Chiang et al. (2023) 13B 0.26 0.18 0.14 0.46 0.28
Qwen-14B-Chat Bai et al. (2023) 14B 0.30 0.24 0.21 0.46 0.29
WizardLM-13B-v1.2 Xu et al. (2024) 13B 0.02 0.04 0.00 0.00 0.00
Few Flan-T5-Large Chung et al. (2022) 780M 0.16 0.11 0.07 0.29 0.17
Flan-T5-XL Chung et al. (2022) 3B 0.16 0.09 0.00 0.29 0.19
Flan-T5-XXL Chung et al. (2022) 11B 0.21 0.14 0.07 0.40 0.21
ChatGLM2-6B Du et al. (2022a) 6B 0.01 0.01 0.00 0.00 0.00
Baichuan2-7B-Chat Baichuan (2023) 7B 0.00 0.00 0.00 0.00 0.00
Baichuan2-13B-Chat Baichuan (2023) 13B 0.02 0.03 0.00 0.06 0.00
InternLM-Chat-7B Team (2023) 7B 0.01 0.00 0.00 0.00 0.01
Vicuna-7B-v1.5 Chiang et al. (2023) 7B 0.01 0.01 0.00 0.00 0.00
Vicuna-13B-v1.5 Chiang et al. (2023) 13B 0.01 0.01 0.00 0.00 0.00
Qwen-14B-Chat Bai et al. (2023) 14B 0.25 0.29 0.07 0.40 0.17
WizardLM-13B-v1.2 Xu et al. (2024) 13B 0.01 0.01 0.00 0.00 0.00
Zero Flan-T5-Large Chung et al. (2022) 780M 0.24 0.29 0.07 0.37 0.15
Flan-T5-XL Chung et al. (2022) 3B 0.25 0.22 0.07 0.43 0.22
Flan-T5-XXL Chung et al. (2022) 11B 0.28 0.29 0.21 0.43 0.21
ChatGLM2-6B Du et al. (2022a) 6B 0.15 0.10 0.00 0.34 0.13
Baichuan2-7B-Chat Baichuan (2023) 7B 0.27 0.25 0.14 0.46 0.22
Baichuan2-13B-Chat Baichuan (2023) 13B 0.29 0.25 0.21 0.46 0.25
InternLM-Chat-7B Team (2023) 7B 0.28 0.28 0.14 0.43 0.22
Vicuna-7B-v1.5 Chiang et al. (2023) 7B 0.22 0.16 0.14 0.46 0.17
Vicuna-13B-v1.5 Chiang et al. (2023) 13B 0.25 0.18 0.21 0.40 0.26
Qwen-14B-Chat Bai et al. (2023) 14B 0.30 0.32 0.21 0.49 0.21
WizardLM-13B-v1.2 Xu et al. (2024) 13B 0.01 0.01 0.00 0.00 0.00
Table 4: Performances of smaller models in three inference-only settings. CoT: The chain-of-thought setting. Few: The few-shot setting. Zero: The zero-shot setting. More details are in Appendix A. NA: Necessary Assumption. SA: Sufficient Assumption. S: Strengthen. W: Weaken.
@@ 5.2 Fine-Tuning with Counterfactual Data
Figure 4: A successful case of counterfactual modification (reasoning type: Sufficient Assumption) generated by GPT-4, which makes a logically consistent ${}^{\prime}$ and is accordant with human reasoning.
@@
Figure 5: An inferior case of counterfactual modification (reasoning type: Weaken) generated by GPT-4. It modifies Statement 1 by replacing intellectual activities with physical activities, where the logical restriction is not satisfied.
5.3 Performances of Small Language Models
We further test various language models on smaller scales and the results are shown in Table 4. All the models in Table 4 directly perform inference without further fine-tuning thus examining their original counterfactual abilities. The detailed settings are demonstrated in Appendix A, and the brief introductions of the language models are listed in Appendix B. We have the following findings: (1) Generally Speaking, all models perform inferiorly. Some of the models, for example, Baichuan2-7B-Chat in the few-shot setting, hardly solve any of the counterfactual questions. Among the models, Qwen-14B-Chat performs the best in both the chain-of-thought setting and the zero-shot setting. But all models perform inferior to large language models or human (2) The models perform better counterfactual reasoning with step-by-step reasoning processes (the CoT setting) while seeing more demonstrations in the prompt (the few-shot setting) harms the performances. It is indicated that the counterfactual cases have obscure reasoning patterns that are challenging for the models to transfer to unseen cases. (3) The performances are correlated to the model scales, but the gaps are not necessarily significant. For instance, Flan-T5-XXL (11B) performs generally better than Flan-T5-Large (780M), and Vicuna-13B-v1.5 performs better than Vicuna-7B-v1.5. (4) For the four different relations, it is shown that Sufficient Assumption and Weaken are significantly more challenging than the other two reasoning types. The possible reason is that compared to the other two reasoning types, Sufficient Assumption and Weaken require more reasoning steps such as reversed thinking. Two cases are further shown in Section 5.4. To sum up, the CLoMo task is challenging to current language models. Therefore, further investigation is needed on the counterfactual reasoning abilities of language models.
@@ 5.4 Case Study
We compare the modification by GPT-4 and humans. Figure 4 shows a case about sufficient assumption. That is, whether premise is a sufficient assumption for argument. The focus of the two premises here is to predict the impact of the charter provision. Switching from Premise 1 to Premise 2, the focus of discussion changes from the group of major powers to other nations in response to threats to world peace. This mainly affects the elaboration of Statement 3. GPT-4 has revised Statement 3 accordingly. Humans also made changes to Statement 3. Both revisions emphasized the influence of the five major powers in Statement 3, which corresponded to Premise 2. We find that GPT-4 can handle complex logical reasoning and counterfactual reasoning to a certain extent.
Figure 5 is about weakening an argument. The argument is on human intellectual development. Statement 2 provides evidence to support Statement 1. Premise 1 is on medical conditions and treatments providing counterexamples, and Premise 2 is on inaccuracies in research data. Human modifies Statement 2 to emphasize conversation, thus satisfying the logical conflict with the inaccuracies in research data described in Premise 2, thus satisfying the weaken relation. GPT-4 modifies Statement 1 by replacing intellectual activities with physical activities such as sports or gym. Intellectual/physical activities have some counterfactual contrast, but in the context of the argument at hand, the logical relation restriction is not satisfied. We find that GPT-4 can still be flawed in counterfactual reasoning. In conclusion, complex counterfactual reasoning is challenging for large language models and needs improvements.